OpenStudy (toxicsugar22):

For the following reaction, 5.30 grams of propane (C3H8) are mixed with excess oxygen gas. The reaction yields 13.1 grams of carbon dioxide. C3H8 (g) + 5 O2 (g) = 3 CO2 (g) + 4 H2O (g)

OpenStudy (anonymous):

1mole C3H8 = 3(12)+8=36+8=44g yields 3 moles of CO2 = 3 ( 12+32 )=3(44)g. (a) 5.3g of C3H8 should yield 3(5.3)=15.9g (b)13.1*100/15.9 = %yield

OpenStudy (toxicsugar22):

For the following reaction, 6.99 grams of hydrogen gas are mixed with excess nitrogen gas. The reaction yields 27.7 grams of ammonia. N2 (g) + 3 H2 = (g) 2 NH3 (g)

OpenStudy (toxicsugar22):

What is the theoretical yield of ammonia ? grams (2) What is the percent yield for this reaction ? %

OpenStudy (toxicsugar22):

can you help me

OpenStudy (toxicsugar22):

please

OpenStudy (anonymous):

follow the same steps

OpenStudy (anonymous):

3 moles of H2 should yield 2 moles of NH3

OpenStudy (anonymous):

expected yield of NH3 = 6.99*2/3